Grade 11Discuss with Askiitians Tutors

page 57 q4 of center of mass module iam class 11 of homia22batch

Arnab Biswas
11 Years agoGrade 11

1 Answer

Shobhit Varshney
11 Years ago
Hi,

In this question lets find out the angle of deflection (a) of two spheres from horizontal after they are being hit by another sphere.
from the geometry, sin a = r/2r, a = 30o

Applying conservation of momentum in x-direction,
mu = mv cos30 + mv cos30, v = u /\sqrt{3}

e = v cos30 / u = ½
You can now easily find out loss in kinetic energy by calculating inital and final kinetic energy in terms of u

thanks